# DISRetrieval: Harnessing Discourse Structure for Long Document Retrieval

### v1

@@ Appendix C Extension of Technical Details
C.1 Discourse Parser Details
@@
$\mathbf{h}_{v}=\begin{cases}\text{PLM}(s_{i}),&\text{if }v\text{ is a sentence}\\ \frac{1}{|C(v)|}\sum_{u\in C(v)}\mathbf{h}_{u},&\text{if }v\text{ is a relationship node}\end{cases}$ (4)
@@
Input: Document $D$ containing sentences $S_{i}$, paragraphs $P_{i}$; Query $q$
Output: Retrieved evidence segments $E$
/* Stage 1: Discourse-Aware Tree Construction */
for each paragraph $P_{i}\in D$ do
       Initialize stack $\sigma$ and sentence queue $\beta\leftarrow S_{i}$;
       while $\beta$ not empty OR $|\sigma|>1$ do
             /* RST parsing operations */
[…]
              // Determine next parsing action
             switch $a_{t}$ do
                   case shift do
                         Move next sentence from $\beta$ to $\sigma$;
                          // Add new sentence to stack
                   end case
                  case reduce do
                         $s_{2},s_{1}\leftarrow$ Pop top two elements from $\sigma$;
                         Determine discourse relation $r$ between $s_{1}$ and $s_{2}$;
                          // Using RST relations
                         Create new node with relation $r$;
                         Push new node to $\sigma$;
                   end case
                  case pop_root do
                         Final tree node $\leftarrow$ Pop from $\sigma$;
                          // Complete the tree
                   end case
             end switch
       end while
      $T_{i}\leftarrow$ Resulting discourse tree;
      /* LLM Enhancement for Paragraph Tree */
       for each non-leaf node $v\in T_{i}$ (bottom-up order) do
             $t_{l},t_{r}\leftarrow$ Get content from left and right children;
             if $|t_{l}|+|t_{r}|\geq\tau$ then
                   $t_{v}\leftarrow f_{\text{LLM}}(t_{l},t_{r})$;
                    // Generate concise summary using LLM
            else
                   $t_{v}\leftarrow t_{l}\oplus t_{r}$;
                    // Direct concatenation for short text
             end if
       end for
      Store root text representation $t^{i}_{\text{root}}$ for $T_{i}$;
end for
Initialize stack $\sigma$ and queue $\beta\leftarrow\{t^{1}_{\text{root}},...,t^{n}_{\text{root}}\}$;
while $\beta$ not empty OR $|\sigma|>1$ do
       /* Similar RST parsing at document level */
       Apply RST parsing operations as in Phase 1;
       Build document tree $T_{\text{doc}}$;
end while
/* Stage 2: Node Representation */
for each non-leaf node $v\in T_{\text{doc}}$ (bottom-up order) do
       $t_{l},t_{r}\leftarrow$ Get content from left and right children;
       if $|t_{l}|+|t_{r}|\geq\tau$ then
             $t_{v}\leftarrow f_{\text{LLM}}(t_{l},t_{r})$;
              // LLM-based enhancement for document-level nodes
      else
             $t_{v}\leftarrow t_{l}\oplus t_{r}$;
              // Direct concatenation for short text
       end if
end for
$T_{D}\leftarrow$ Replace leaf nodes in $T_{\text{doc}}$ with corresponding $T_{i}$;
  // Integrate trees into unified structure
for each node $v\in T_{D}$ do
       $e_{v}\leftarrow\text{Encoder}(t_{v})$;
        // Generate dense vector representations
end for
/* Stage 3: Hierarchical Evidence Retrieval and Selection */
$e_{q}\leftarrow\text{Encoder}(q)$;
  // Transform query to embedding space
$\text{scores}\leftarrow\{\}$, $E\leftarrow\{\}$;
for node $v\in T_{D}$ do
       $\text{scores}[v]\leftarrow\text{cosine}(e_{q},e_{v})$;
        // Compute relevance scores
end for
Apply Algorithm 1 for evidence selection;
return $E$
Algorithm 2 DISRetrieval: Discourse Structure-based Long Document Retrieval
C.2 Iterative Tree Construction Process
@@
    • For each non-leaf node $v$ with children $c_{l}$ and $c_{r}$:
      $t_{v}=\begin{cases}f_{\text{LLM}}(t_{l},t_{r}),&\text{if }|t_{l}|+|t_{r}|\geqslant\tau\\ t_{l}\oplus t_{r},&\text{otherwise}\end{cases}$ (8)
      where $t_{l}$ and $t_{r}$ are the textual content of child nodes
@@
    • For each non-leaf node $v\in T_{\text{doc}}$ with children $c_{l}$ and $c_{r}$:
@@
Appendix D Detailed Experiment Settings
D.1 Dataset Specifications
@@
The QuALITY dataset consists of long passages primarily drawn from fiction stories and magazine articles, with an average length significantly longer than typical QA datasets. Each question is accompanied by multiple choice options that test comprehensive understanding of the passage. We utilize the validation set with publicly available labels for our experiments to enable thorough analysis and comparison. Notably, to enable more extensive experimentation, we evaluated on the validation set with publicly available labels rather than the submission-required test set. More detailed information is presented in Table 4.
@@
Dataset Used Set Question Num Document Num Avg. Document Length (words)
QASPER test 1456 416 4170
QuALITY dev 2086 115 5022
D.2 Model Specifications
@@ D.3 Implementation Details
D.4 Evaluation Protocol
For QASPER’s retrieval evaluation, we compute token-level F1 and Recall scores by comparing the retrieved context against the annotated ground truth evidence spans. This provides a fine-grained assessment of retrieval quality beyond simple overlap metrics. For QuALITY, the accuracy metric reflects the proportion of correctly answered multiple-choice questions, directly measuring the impact of retrieval quality on downstream QA performance.
D.5 Experiments Compute Resources
@@ D.6 Prompts
Answer the following multiplie-choice question:{question}
@@
Appendix E Extended Experiment Results and Analyses
E.1 Analysis of Multi-granularity Adaptive Retrieval
@@
Table 5: Statistical analysis of retrieved intermediate node characteristics across QuALITY and QASPER datasets. The table compares various key metrics: Avg. Sentence Length means the average sentence length of all documents; Avg. Mid Node Depth is the average depth of retrieved intermediate nodes; Avg. Leaf Num represents the average number of leaf nodes that each retrieved intermediate node maps to; Avg. Mid Node Percentage is the average percentage of intermediate nodes among the Top-20 retrieved nodes.
Dataset Avg. Sentence Length Avg. Mid Node Depth Avg. Leaf Num Avg. Mid Node Percentage / %
SBERT OpenAI SBERT OpenAI SBERT OpenAI
QASPER 22.32 5.29 5.53 12.64 13.37 54.15 65.13
QuALITY 14.86 13.78 17.09 67.67 88.67 73.93 88.53
@@
Figure 7: Ablation results different values of $K$. The horizontal axis represents different choices of $K$, and the vertical axis indicates generation performance (F1-match for QASPER and accuracy for QuALITY). All question answering tasks are conducted on the UnifiedQA-3B model.
E.2 Ablation study on different Top-K settings
@@
E.3 Detailed Results of different answer types on QASPER Dataset
The question-answering task on the QASPER dataset includes four types of answers: Abstractive, Yes / No, Extractive and Answerable / Unanswerable.
  • Abstractive: This type of answers use abstract and summarized content rather than the original text from the document. They often involve rephrasing or synthesizing content to provide a concise and coherent response.
  • Yes / No: These answers are straightforward and binary, simply Yes or No to the question.
  • Extractive: These answers are directly taken from the original text. The goal is to pinpoint and extract the specific segment of text that directly answers the question.
  • Answerable / Unanswerable: This type categorizes whether the question can be answered based on the information provided in the text. These type of answers correspond to questions that the original documents do not contain sufficient supporting information, and the model is expected to directly reply with "unanswerable."
We present the detailed results for each type of answer in Table 6. Notably, both GPT-4.1-mini and Deepseek-v3 achieve zero scores on the Answerable / Unanswerable type due to their inability to meet the specific answer word requirement.
Table 6: Question answering performance across different answer types on QASPER dataset. Abs, Bool, Ext, and None represent Abstractive, Yes / No, Extractive and Answerable / Unanswerable answers respectively. Final denotes the average F1 score across different answer types.
Method 200 300 400
Abs Bool Ext None Final Abs Bool Ext None Final Abs Bool Ext None Final
UnifiedQA-3B + SBERT
flatten-chunk 16.37 73.30 32.68 18.85 33.97 16.26 75.45 35.16 18.18 35.41 15.73 75.91 36.82 15.83 36.13
flatten-sentence 16.69 75.34 34.46 17.74 35.16 16.69 76.68 38.20 15.13 37.24 17.61 77.48 39.6 11.67 37.99
raptor 16.26 71.95 32.59 18.25 33.57 16.32 72.48 35.19 17.36 34.95 16.97 77.27 37.28 18.03 37.00
Ours
   Bisection 16.13 77.38 36.53 15.57 36.17 16.57 76.92 39.49 12.71 37.66 17.03 75.78 41.81 10.34 38.84
   DISRetrieval 17.30 77.48 37.63 15.57 37.26 18.55 75.68 41.02 11.57 38.83 18.50 77.48 42.96 13.33 40.25
UnifiedQA-3B + OpenAI Embedding
flatten-chunk 16.33 75.45 38.77 18.03 37.50 17.39 76.13 39.95 17.24 38.46 17.68 77.27 41.06 17.36 39.03
flatten-sentence 17.33 75.45 40.86 13.56 38.43 17.53 76.71 42.11 13.68 39.31 18.69 77.06 43.16 10.43 40.06
raptor 15.68 74.21 39.51 15.65 37.46 16.91 77.03 42.72 15.18 39.96 16.47 76.92 42.39 13.16 39.53
Ours
   Bisection 17.27 73.76 39.70 11.11 37.41 18.03 76.47 42.77 8.33 39.49 18.21 77.73 43.22 5.88 39.70
   DISRetrieval 18.13 77.73 41.5 9.24 38.91 19.67 77.93 43.12 7.69 40.24 20.02 77.38 44.57 5.13 40.80
GPT-4.1-mini + SBERT
flatten-chunk 21.16 78.09 38.48 37.37 23.85 84.55 42.78 41.38 24.90 84.17 44.76 42.72
flatten-sentence 22.06 79.11 42.19 39.82 23.46 80.93 44.78 41.84 23.80 78.42 46.29 42.44
raptor 20.62 82.13 37.70 37.55 22.42 82.21 41.45 39.95 24.73 83.57 44.46 42.50
Ours
   Bisection 22.61 84.13 42.58 40.98 23.54 82.53 45.87 42.74 24.44 80.01 47.12 43.49
   DISRetrieval 22.40 79.54 44.30 41.04 24.19 83.03 45.29 42.58 25.20 82.51 48.51 44.52
GPT-4.1-mini + OpenAI Embedding
flatten-chunk 24.12 79.46 43.19 41.13 26.90 83.09 45.45 43.34 27.90 80.65 48.29 44.78
flatten-sentence 25.34 79.35 46.75 43.08 27.13 82.19 49.12 45.28 26.94 83.30 49.60 45.78
raptor 23.54 78.85 43.36 40.88 26.08 81.21 45.95 43.26 25.94 80.59 47.40 43.85
Ours
   Bisection 24.99 81.78 46.12 43.12 25.37 82.60 48.41 44.54 27.81 80.48 49.97 45.69
   DISRetrieval 25.04 83.57 46.68 43.47 26.96 82.82 49.45 45.47 27.14 82.52 50.43 45.97
Deepseek-v3 + SBERT
flatten-chunk 17.23 81.70 35.27 35.29 17.50 85.65 38.86 37.99 20.11 85.71 42.48 40.57
flatten-sentence 17.74 80.26 36.99 36.51 18.58 81.86 41.61 39.41 20.12 83.04 43.45 40.82
raptor 15.75 80.27 35.03 34.66 17.76 82.06 38.53 37.30 19.73 83.86 41.66 39.77
Ours
   Bisection 18.05 78.76 38.83 37.28 18.24 85.40 41.76 39.80 19.60 82.38 44.10 40.96
   DISRetrieval 17.18 83.93 38.22 37.44 18.10 85.40 42.92 40.28 19.65 85.71 44.70 41.86
Deepseek-v3 + OpenAI Embedding
flatten-chunk 19.20 82.14 39.06 38.14 21.81 82.96 42.75 40.78 21.77 84.62 44.90 42.26
flatten-sentence 20.45 81.70 41.02 39.27 22.03 82.06 44.73 41.57 23.23 84.68 45.70 43.04
raptor 17.47 80.89 39.55 37.84 20.63 82.88 43.10 40.50 21.56 84.82 44.74 42.17
Ours
   Bisection 20.62 81.86 40.73 39.27 22.20 84.51 44.15 41.81 23.13 84.89 47.09 43.57
   DISRetrieval 20.82 81.61 42.13 40.01 21.95 83.93 44.73 42.10 23.06 84.96 45.90 43.21

### v2

@@ Appendix A Extension of Technical Details
A.1 Discourse Parser Details
@@
$\mathbf{h}_{v}=\begin{cases}\text{PLM}(s_{i}),&\text{if }v\text{ is {sentence}}\\ \frac{1}{|C(v)|}\sum_{u\in C(v)}\mathbf{h}_{u},&\text{if }v\text{ is {relationship}}\end{cases}$ (7)
@@
A.2 Iterative Tree Construction Process
@@
    • For each non-leaf node $v$ with children $v_{l}$ and $v_{r}$:
      $v^{*}=\begin{cases}f_{\text{LLM}}(v_{l},v_{r}),&\text{if }|v_{l}|+|v_{r}|\geqslant\tau\\ f_{\text{merge}}(v),&\text{otherwise}\end{cases}$ (11)
      where $|v_{l}|$ and $|v_{r}|$ are the textual content length of child nodes
@@
    • For each non-leaf node $v\in T_{\text{doc}}$ with children $v_{l}$ and $v_{r}$:
@@
A.3 Detailed LLM Enhancement Process
This section provides a detailed illustration of the bottom-up LLM enhancement process described in Phase 2 of our discourse tree construction methodology (Section 3.1.2).
A.3.1 LLM Enhancement Workflow
Figure 5 demonstrates the complete workflow of our LLM enhancement process using a concrete example from a research paper paragraph. The process consists of three main components:
Input Structure.
The top portion shows the input paragraph with its initial sentence-level discourse tree structure. Each sentence is represented as a leaf node, connected through discourse relations such as Contrast, List, Summary, and Topic. This hierarchical structure captures the rhetorical organization of the paragraph content.
Adaptive Processing Strategy.
The center portion illustrates our two-way processing strategy based on text length thresholds. When the combined length of child nodes ($|t_{l}|+|t_{r}|$) falls below the threshold $\tau$, we apply direct concatenation ($\bigoplus$) to preserve the original textual details. Conversely, when the combined length exceeds $\tau$, we employ LLM-based summarization to generate more concise representations while retaining essential semantic information.
Enhanced Tree Structure.
The bottom portion shows the resulting enhanced discourse tree with progressively generated semantic summaries. Each internal node now contains meaningful textual representations that capture the essence of its subtree content. For example, the Contrast relation between different learning methods is transformed into the concrete summary "Different learning methods suit different applications", while maintaining the hierarchical discourse structure.
A.3.2 Key Benefits
This LLM enhancement process provides several advantages:
  • Semantic Preservation: Maintains essential meaning while reducing textual complexity
  • Computational Efficiency: Adaptive thresholding prevents unnecessary LLM calls for short text segments
  • Hierarchical Coherence: Preserves discourse relationships at multiple granularity levels
  • Downstream Compatibility: Generates representations suitable for sentence-level discourse parsing
The transformation from abstract rhetorical relations to concrete semantic representations enables our framework to effectively bridge the gap between discourse structure and semantic understanding, facilitating improved retrieval performance in long document question answering tasks.
Appendix B Detailed Experiment Settings
B.1 Dataset Specifications
@@
Dataset Used Set Question Num Doc. Num Avg. words Max. words
QASPER test 1456 416 4170 21,165
QuALITY dev 2086 115 5022 5967
NarrativeQA test 10,577 355 51,372 346,902
@@
The QuALITY dataset consists of long passages primarily drawn from fiction stories and magazine articles, with an average length longer than typical QA datasets. Each question is accompanied by multiple choice options that test comprehensive understanding of the passage. We utilize the validation set with publicly available labels for our experiments to enable thorough analysis and comparison. Notably, to enable more extensive experimentation, we evaluated on the validation set with publicly available labels rather than the submission-required test set.
The NarrativeQA dataset is a significant advancement in evaluating machine reading comprehension. Unlike traditional datasets that focus on short passages, NarrativeQA emphasizes understanding long-form narratives, such as entire books or full-length movie scripts. One of the defining features of NarrativeQA is the length of its documents. On average, each document contains 51,372 words. Notably, some individual documents extend up to 346,902 words, making it one of the longest reading comprehension datasets available. More detailed information is presented in Table 6.
B.2 Model Specifications
@@ B.3 Implementation Details
B.4 Experiments Compute Resources
@@ B.5 Prompts
Answer the following multiple-choice question:{question}
@@
Appendix C Extended Experiment Analysis
C.1 Computational Efficiency and Scalability
To evaluate the computational efficiency and scalability of our approach, we conducted comprehensive timing analysis across varying document lengths. We systematically selected documents ranging from 10,000 to 100,000 words, performed discourse-aware tree construction and LLM enhancement, and measured the computation time for each process. All experiments were conducted on a single A800 80G GPU, with LLM enhancement performed using the LLaMA3.1-8B-Instruct model accelerated by the vLLM framework. To ensure statistical reliability, we sampled 10 documents for each length range and report the average processing time with standard deviation. The detailed results are presented in Figure 6.
Figure 6: Average processing time of discourse-aware tree construction and LLM enhancement across varying document lengths.
The processing time for both discourse-aware tree construction and LLM enhancement exhibits near-linear scaling with document length, demonstrating favorable computational complexity for long documents. Specifically, tree construction for documents up to 100,000 words requires less than 40 seconds, while LLM enhancement completes within 140 seconds, confirming practical feasibility for real-world applications. The linear scaling behavior indicates that our method maintains consistent per-word processing efficiency regardless of document length. Moreover, as discussed in Section 4.3, the size and performance of the summarization model have no significant impact on the results, which further suggests that our method can be deployed with a modest computational cost.
Crucially, both tree construction and LLM enhancement are query-independent preprocessing operations performed once per document, allowing costs to be amortized across multiple queries on the same document. For instance, processing a 50,000-word document requires approximately 90 seconds of preprocessing but can subsequently serve unlimited queries with minimal additional overhead. In practical deployment scenarios where documents receive multiple queries over time, this one-time preprocessing cost becomes negligible compared to the cumulative benefits of improved retrieval accuracy.
Compared to traditional chunking approaches that require no preprocessing, our method introduces modest upfront costs but delivers significantly better retrieval quality. The preprocessing overhead is comparable to other hierarchical methods like RAPTOR, while providing superior performance through linguistically-grounded discourse modeling.
C.2 Analysis of Multi-granularity Adaptive Retrieval
@@
Figure 8: Ablation results with different values of $K$. The horizontal axis represents different choices of $K$, and the vertical axis indicates generation performance (F1-match for QASPER and accuracy for QuALITY). All question answering tasks are conducted on the UnifiedQA-3B model.
C.3 Ablation study on different Top-K settings
@@
Figure 9: Ablation results with different values of $\tau$ across three datasets. Sentence-BERT is used as the embedding model and UnifiedQA-3B as the question-answering model, with a context length of 400 words.
C.4 Ablation study on parameter $\tau$ across different datasets
Equation 2 employs a threshold parameter $\tau$ to decide whether use an LLM for summarization or directly merge the subtree. To assess the impact of this parameter and identify the optimal settings, we conduct experiments on three datasets. The results are shown in figure 9.
Performance Trends.
For QASPER, the F1 score is highest when $\tau=0$ and gradually declines as $\tau$ increases, indicating that summarization is more beneficial than direct merging sentences of subtrees. In contrast, QuALITY and NarrativeQA reach their peak performance at $\tau=50$ and gradually decline afterward, only approaching the level observed at $\tau=0$ when $\tau$ reaches 300 and 200 respectively, suggesting that a moderate merging strategy is essential for these datasets.
Dataset Analysis
These results are consistent with the characteristics of the datasets. QASPER consists of academic papers, where sentences are relatively independent and logically complete. So keeping sentences as the minimal unit allows for more precise information extraction. In contrast, QuALITY and NarrativeQA mainly contain novels and movie scripts, which include numerous dialogues and short sentences that require a relatively complete context to convey the correct semantics. For these two datasets, merging subtrees based on rhetorical structure parsing results enables logically related sentences to form more coherent text segments, preventing semantic fragmentation. At the same time, moderate merging is crucial, as excessively long text segments may introduce redundancy. Therefore, we set $\tau=0$ for QASPER, $\tau=50$ for QuALITY and NarrativeQA respectively. This can also provide guidance for other datasets: for documents with more formal and complete sentences, smaller $\tau$ values are preferable, whereas for documents with shorter and more fragmented sentences, relatively larger $\tau$ values should be used.